OpenStudy (anonymous):

Simplify: 16 – 4 + 14+ 2 / I keep geting 13, but it is telling me its wrong. People say 28. But my options are only 13. 20. 19. 1.HELP

OpenStudy (anonymous):

can you explain how you got 13 ?

OpenStudy (anonymous):

yes please.

OpenStudy (anonymous):

(16-4+14)/2=13

OpenStudy (anonymous):

but the way you posted the question there are only addition and subtraction operations.

OpenStudy (anonymous):

good job dude :D

OpenStudy (anonymous):

Freak. i ment to put a divide sign instead of plus. it should be 14\2. i went from left to right. divided 14/2 got 7. -4+7=3. 16-3. but then i relized you cant put the minus sign as a negative and crap then did 4+7=11 instead and 16-11=5. amd thts not my option either. Thats what the instructions sayyyyy to do

OpenStudy (anonymous):

So the question is this ? 16 – 4 + 14/ 2

OpenStudy (anonymous):

if so, the answer is 19 not 13.

OpenStudy (anonymous):

you need to do multiplication and division operations before addition and substraction

OpenStudy (anonymous):

so, we start with \[16-4+14/2\] first we do the division, and get this: \[16-4+7\] And now we do the additions and subtractions, and get \[19\]